### v1

# From Automation to Autonomy: A Survey on Large Language Models in Scientific Discovery

Abstract
Large Language Models (LLMs) are catalyzing a paradigm shift in scientific discovery, evolving from task-specific automation tools into increasingly autonomous agents and fundamentally redefining research processes and human-AI collaboration. This survey systematically charts this burgeoning field, placing a central focus on the changing roles and escalating capabilities of LLMs in science. Through the lens of the scientific method, we introduce a foundational three-level taxonomy—Tool, Analyst, and Scientist—to delineate their escalating autonomy and evolving responsibilities within the research lifecycle. We further identify pivotal challenges and future research trajectories such as robotic automation, self-improvement, and ethical governance. Overall, this survey provides a conceptual architecture and strategic foresight to navigate and shape the future of AI-driven scientific discovery, fostering both rapid innovation and responsible advancement.
[…]
tcb@breakable
From Automation to Autonomy:
A Survey on Large Language Models in Scientific Discovery
Tianshi Zheng, Zheye Deng, Hong Ting Tsang, Weiqi Wang, Jiaxin Bai, Zihao Wang, Yangqiu Song Department of Computer Science and Engineering, HKUST, Hong Kong SAR, China https://github.com/HKUST-KnowComp/Awesome-LLM-Scientific-Discovery
1 Introduction
The relentless advancement of Large Language Models (LLMs) has unlocked a suite of emergent abilities, such as planning Huang et al. (2024b), complex reasoning Huang and Chang (2023), and instruction following Qin et al. (2024). Moreover, integrating agentic workflows enables LLM-based systems to perform advanced functions, including web navigation He et al. (2024), tool use Qu et al. (2025), code execution Jiang et al. (2024a), and data analytics Sun et al. (2024). In scientific discovery, this convergence of advanced LLM capabilities and agentic functionalities is catalyzing a significant paradigm shift. This shift is poised not only to accelerate the research lifecycle but also to fundamentally alter the collaborative dynamics between human researchers and artificial intelligence in the pursuit of knowledge.
Figure 1: Stages of the scientific method with corresponding LLM applications and research topics.
Autonomy Levels LLMs’ Role Human’s Role Task Scope Agentic Workflow
[…]
LLM as Tool
Task Automation Tool Task Allocation Explicitly Defined Simple & Static
[…]
LLM as Analyst
Data Modeling &
Analytical Agent
Problem Definition &
Output Validation
Goal-Oriented Advanced
[…]
LLM as Scientist
Open Exploratory &
Discovery Agent
Minimal Intervention Open-Ended Strategic & Iterative
Table 1: Three levels of autonomy in LLM-based scientific discovery.
However, this rapid expansion of LLM applications and the ongoing paradigm shift in scientific discovery present notable challenges. The accelerated pace of LLM evolution and their deepening integration into complex research complicate systematic assessment, necessitating conceptual frameworks to structure current understanding and chart future directions. While existing surveys have provided valuable overviews of LLMs in various scientific domains Zhang et al. (2024, 2025) or have cataloged particular AI techniques for science Luo et al. (2025); Reddy and Shojaee (2025), they often focus on discipline-specific applications or a static snapshot of LLM capabilities. Consequently, existing reviews may overlook the crucial trend of increasing LLM autonomy and their evolving roles across the entire scientific method, leaving their comprehensive impact and trajectory towards greater independence underexplored.
To systematically chart this evolving landscape and address the identified gap, we anchor our analysis in the six stages (Figure 1) of the established scientific method Popper (1935); Kuhn (1962): (1) observation and problem definition, (2) hypothesis development, (3) experimentation and data collection, (4) data analysis and interpretation, (5) drawing conclusions, and (6) iteration and refinement. Our examination of LLM applications across these stages reveals a significant trend: LLMs are progressing from performing discrete, task-oriented functions within a single stage to deployment in sophisticated, multi-stage agentic workflows. Notably, emerging research Schmidgall et al. (2025); Yamada et al. (2025) now explores developing LLM-based systems capable of autonomously navigating nearly all these stages. To effectively capture and delineate this trajectory of increasing capability and independence, we introduce a foundational three-level taxonomy for LLM involvement in scientific discovery (Table 1): (i) LLM as Tool, where models augment human researchers by performing specific, well-defined tasks under direct supervision; (ii) LLM as Analyst, where models exhibit greater autonomy in processing complex information, conducting analyses, and offering insights with reduced human intervention; and (iii) LLM as Scientist, representing a more advanced stage where LLM-based systems can autonomously conduct major research stages, from formulating hypotheses to interpreting results and suggesting new avenues of inquiry.
Building upon this taxonomic framework, we further identify critical gaps in the current research landscape and highlight pivotal challenges and future trajectories for the field, including: (1) fully autonomous discovery cycles for evolving scientific inquiry without human intervention; (2) robotic automation for interaction in the physical world for laboratory experimentation; (3) continuous self-improvement and adaptation from past research experiences; (4) transparency and interpretability of LLM-conducted research; and (5) ethical governance and societal alignment. Addressing these multifaceted challenges will be crucial for achieving a future where AI acts as a transformative partner in scientific exploration.
This survey focuses on LLM-based systems in scientific discovery, particularly their varying levels of autonomy. While acknowledging the broad impact of LLMs in science, we deliberately narrow our scope to exclude research on general-purpose scientific LLMs or LLMs for domain-specific scientific knowledge acquisition and reasoning, which are well covered in existing surveys Zhang et al. (2024, 2025). The remainder of this paper is organized as follows: Section 2 details our taxonomy and its interaction with the scientific method. Section 3 presents LLM as Tool applications, categorized by scientific method stages. Section 4 examines LLM as Analyst works by scientific domain, while Section 5 analyzes LLM as Scientist systems, focusing on their idea development and refinement strategies. Section 6 explores challenges and future directions.
Figure 2: Taxonomy of research works in LLM-based scientific discovery with detailed categorization.
2 Three Levels of Autonomy
Table 1 illustrates the three levels of autonomy in LLM-based scientific discovery with their associated features. In this section, we discuss their applications and characteristics in more detail.
LLM as Tool (Level 1). Level 1 represents the most foundational application of LLMs in scientific discovery. At this stage, LLMs function primarily as tailored tools under direct human supervision, designed to execute specific, well-defined tasks within a single stage of the scientific method. Their role is to augment human capabilities by automating or accelerating discrete activities such as literature summarization, drafting initial text for manuscripts, generating code snippets for data processing, or reformatting datasets. The autonomy of LLMs at this level is limited; they operate based on explicit human prompts and instructions, with outputs typically requiring human validation and integration into the broader research workflow. The primary goal is to enhance researcher efficiency and reduce routine task burdens.
LLM as Analyst (Level 2). In Level 2, LLMs exhibit a greater degree of autonomy and move beyond purely static, task-oriented applications. Here, LLMs function as passive agents, capable of more complex information processing, data modeling, and analytical reasoning with reduced human intervention for intermediate steps. While still operating within boundaries set by human researchers, these systems can independently manage sequences of tasks, such as analyzing experimental datasets to identify trends, interpreting outputs from complex simulations, or even performing iterative refinement of models. The human researcher typically defines the overall analytical goals, provides the necessary data, and critically evaluates the insights or interpretations generated by the LLM.
LLM as Scientist (Level 3). Level 3 applications signify a significant leap in autonomy, where LLM-based systems operate as active agents capable of orchestrating and navigating multiple stages of the scientific discovery process with considerable independence. These systems can demonstrate initiative in formulating hypotheses, planning and executing experiments, analyzing the resultant data, drawing preliminary conclusions, and potentially proposing subsequent research questions or avenues for exploration. LLM-based systems at this level can drive substantial portions of the research cycle, conducting scientific discovery with minimal human intervention.
Collectively, we present our full taxonomy with detailed categorization in Figure 2, which consolidates 90 research works within our focused scope across all three levels of autonomy.
3 Level 1. LLM as Tool (Table A1)
In this section, we introduce Level 1 research works in LLM-based scientific discovery, categorized by the stages in the scientific method they address.
3.1 Literature Review and Information Gathering
Literature Review
Automatic literature search and retrieval is crucial for identifying research gaps and formulating research questions. Lála et al. (2023) first introduced the literature retrieval benchmark LitQA, featuring a RAG-based agent, PaperQA. LitLLM Agarwal et al. (2024) further provided a comprehensive RAG-based toolkit for LLM-driven literature review. Dennstädt et al. (2024) directed their focus to the biomedical domain, highlighting the potential of LLMs in literature screening. Other approaches, such as SCIMON Wang et al. (2024a) and ResearchAgent Baek et al. (2025), have integrated active literature retrieval with the generation of research ideas. More recently, Gao et al. (2025) tackled the task of hierarchically organizing scientific literature through fine-grained paper abstraction. Nevertheless, several ‘Deep Research’ products OpenAI (2025); Google (2025); xAI (2025) have recently been released, featuring enhanced agentic workflows that support automated literature web search, organization, and report generation, thereby significantly accelerating traditional, human-intensive literature research processes.
Information Aggregation
In parallel, several works have explored effective methods for aggregating information from scientific papers into tabular summaries. ArxivDIGESTables Newman et al. (2024) investigated cross-literature table generation using LLMs, accompanied by an automated evaluation strategy. ArXiv2Table Wang et al. (2025) revised the evaluation protocol and provided a comprehensive benchmark. Methodologies such as Text-Tuple-Table Deng et al. (2024b) and TKGT Jiang et al. (2024b) have enhanced the quality of LLM-based table generation by incorporating tuple-based structures and graph modalities.
3.2 Idea Generation and Hypothesis Formulation
Idea Generation
Numerous research efforts have focused on the automated generation of novel research ideas and hypotheses. In the general domain, benchmarks such as IdeaBench Guo et al. (2024a) and LiveIdeaBench Ruan et al. (2025) have evaluated the capability of LLMs to generate research ideas based on provided literature summaries. Concurrently, LLM-based agent frameworks, including Nova Hu et al. (2024a), SciAgents Ghafarollahi and Buehler (2024b), and KG-CoI Xiong et al. (2024), have been proposed to enhance idea generation through effective reasoning over academic knowledge graphs, iterative planning, and searching. Moreover, several exploratory studies have assessed the novelty and quality of LLM-generated ideas for AI research, underscoring the potential for automated idea generation when coupled with appropriate human guidance Si et al. (2024); Feng et al. (2025); Qiu et al. (2025). Furthermore, many studies within natural science disciplines have investigated LLM-based idea generation in domain-specific contexts. For example, Ciucă et al. (2023) proposed adopting adversarial prompting for effective idea generation in astronomy. In biology, Buehler (2024) enhanced idea generation by integrating knowledge extraction and graph representations.
Hypothesis Formulation
Building upon identified ideas, the design of testable scientific hypotheses has also been a significant focus. Qi et al. (2023) and Yang et al. (2024) examined the ability of LLMs to propose hypotheses, demonstrating their considerable capacity for generating novel yet valid hypotheses under open-ended constraints. Methodologies such as Scideator Radensky et al. (2025) have been developed to investigate human-LLM collaboration to facilitate grounded idea and hypothesis generation. Within natural science, benchmarks Qi et al. (2024) and methods O’Brien et al. (2024) have extended hypothesis generation into the biomedical domain. Meanwhile, MOOSE-Chem Yang et al. (2025) offers a systematic evaluation benchmark and an agent framework specifically for hypothesis discovery in chemistry.
3.3 Experiment Planning and Execution
Experiment planning and execution constitute a crucial stage in LLM-based scientific discovery. While integral to advanced Level 2 and Level 3 agents, this subsection focuses on Level 1 research, where LLMs serve as tools for experimental tasks.
Planning
Regarding experiment planning, Li et al. (2025) discussed the effectiveness of incorporating LLMs into the design of causal discovery experiments. BioPlanner O’Donoghue et al. (2023) introduced an automated evaluation framework for assessing LLMs in biological protocol planning. Furthermore, Shi et al. (2025) proposed a hierarchically encapsulated representation to complement LLMs in biological protocol design.
Execution
For experiment execution, current research has primarily concentrated on code generation, particularly for artificial intelligence research, given the inherent compatibility of terminal interfaces with LLM experimental environments. Early code generation benchmarks, such as ARCADE Yin et al. (2022) and DS-1000 Lai et al. (2022), focused on data science tasks. Subsequent works, including MLE-Bench Chan et al. (2025) and SciCode Tian et al. (2024), incorporate more realistic scenarios, such as those encountered in machine learning engineering and natural science research, thereby presenting significant challenges for LLMs. To address these challenges, AIDE (Jiang et al., 2025) proposed enhancing complex code generation capabilities by adopting tree-search methodologies for code optimization.
3.4 Data Analysis and Organization
Tabular Data
In this stage, LLMs assist the scientific workflow by automating processes related to data organization, presentation, and analysis. For data presented in tabular format, Chain-of-Table Wang et al. (2024c) proposes a method to enhance tabular understanding by incorporating evolving tables within the reasoning chain of LLMs. Concurrently, Deng et al. (2024a) highlight the potential of integrating visual information to improve multimodal understanding, thereby aiding tabular comprehension. More recently, Wu et al. (2025) introduced TableBench, a comprehensive benchmark for table-based question answering under practical industrial scenarios.
Chart Data
Beyond tabular data, charts represent another important format for organizing and storing information derived from experimental data. Early benchmarks, exemplified by ChartQA Masry et al. (2022), examined the capabilities of vision transformers in chart-based question answering. Subsequent works, including CharXiv Wang et al. (2024d) and ChartX Xia et al. (2025), have expanded the scope of chart understanding scenarios by utilizing human-curated chart generation or by incorporating real-world chart data sourced from arXiv preprints. Regarding chart generation, AutomaTikZ Belouadi et al. (2024) formulates the process as TikZ code generation from caption text and has demonstrated the efficacy of fine-tuning LLMs using open scientific figure data. More recently, Text2Chart31 Zadeh et al. (2025) employed reinforcement learning with automated feedback to refine chart generation capabilities within the Matplotlib library.
3.5 Conclusion and Hypothesis Validation
In the concluding stages of research, LLMs can provide feedback on, or verify, claims and conclusions derived from experiments.
Paper Review
In this context, a significant focus of contemporary research involves investigating the utility of LLMs as reviewers for artificial intelligence papers. ReviewerGPT Liu and Shah (2023) initially explored the capability of LLMs to identify deliberately inserted errors within research papers, highlighting the necessity for more robust systems to conduct comprehensive reviews. Zhou et al. (2024a) further evaluated static LLMs in the context of reviewing real-world conference papers using a multiple-choice format. Du et al. (2024) conducted a comprehensive analysis of LLM review quality through extensive human studies and comparisons, revealing weaknesses in their ability to identify deficiencies. ClaimCheck Ou et al. (2025) further investigated the capabilities of LLMs in critiquing research claims, demonstrating that this task remains challenging even for highly advanced models such as OpenAI’s o1 OpenAI (2024). Concurrently, research highlights the potential to address these limitations by incorporating multi-agent systems with specialized roles Tyser et al. (2024); Xu et al. (2025a) or through LLM alignment via reinforcement learning Weng et al. (2025).
Hypothesis Validation
Another important application at this stage is the automatic validation of hypotheses by LLMs. Takagi et al. (2023) demonstrated that LLMs possess considerable capabilities in automatically generating code to verify research hypotheses within simplified machine learning problems. SciReplicate-Bench Xiang et al. (2025) further extended this concept to the replication of real-world research papers. Furthermore, Xu et al. (2025c) have navigated this domain into physics research, aiming to enhance the interpretability of the discovery process through the use of multi-agent workflows.
3.6 Iteration and Refinement
The iterative refinement of research hypotheses, as a distinct area of investigation, has received comparatively less attention in current research. Explanation-Refiner Quan et al. (2024) employed theorem provers to verify and subsequently refine LLM-generated hypotheses. Chain-of-Idea Li et al. (2024a) introduced an LLM-based agent framework designed to organize literature and develop research ideas by building upon or extending existing lines of inquiry. More recently, MC-NEST Rabby et al. (2025) adopted Monte-Carlo Tree Search to iteratively verify and refine scientific hypotheses across multiple research domains.
4 Level 2: LLM as Analyst (Table A2)
In this section, we introduce Level 2 research works in LLM-based scientific discovery, categorized according to their task nature and domains.
4.1 Machine Learning Research
Automated Machine Learning (AutoML) Shen et al. (2024) endeavors to generate high-performing modeling configurations for a given task in a data-driven manner. With the advent of LLM-based agents, several studies have explored their application in the automated modeling of machine learning (ML) tasks. MLAgentBench Huang et al. (2024a) evaluates the capabilities of LLMs in designing and executing ML experiments, revealing that LLM performance is often contingent upon task familiarity derived from their training data. The IMPROVE framework Xue et al. (2025) highlighted the significance of iterative refinement mechanisms within LLM agents for enhancing modeling outcomes. CodeScientist Jansen et al. (2025) incorporated an ML modeling agent with machine-generated ideas, thereby progressing towards more autonomous ML research aimed at performance enhancement. In a related development, BudgetMLAgent Gandhi et al. (2025) leverages curated expert collaboration frameworks to achieve superior results using considerably more cost-effective open-source models.
4.2 Data Modeling and Analysis
Automated data-driven analysis, encompassing statistical data modeling and hypothesis validation, represents a foundational application area for LLM-assisted scientific discovery. InfiAgent-DABench Hu et al. (2024b) benchmarked the capabilities of LLMs in static code generation and execution for data analysis using CSV files. Subsequent benchmarks, such as BLADE Gu et al. (2024) and DiscoveryBench Majumder et al. (2024), have improved evaluation robustness by incorporating real-world research papers and expert-curated analytics within their benchmarking environments. These studies indicate that most LLMs struggle with complex data analytics tasks, even when operating within an agent framework Zheng et al. (2023). To address these challenges, DS-Agent Guo et al. (2024b) proposes to enhance LLM performance by incorporating a case-based reasoning method to improve domain knowledge acquisition. In a related effort, DAgent Xu et al. (2025b) extended the application domain to querying relational databases and enabled report generation using results derived from decomposed sub-problems.
4.3 Function Discovery
Function discovery, which aims to identify the underlying equations from observational data of variables, has been significantly influenced by the advancement of AI-driven symbolic regression (SR) Udrescu and Tegmark (2020); Kamienny et al. (2022). LLM-SR Shojaee et al. (2025a) enhanced function discovery by leveraging the prior domain knowledge of LLMs and incorporating feedback from clustered memory storage. LLM-SRBench Shojaee et al. (2025b) introduced a benchmark for evaluating LLMs as function discovery agents, which incorporates function transformations to mitigate data contamination. Furthermore, other studies have explored the capabilities of LLMs in discovering complex models within specific domains, such as Physics Koblischke et al. (2025) and Statistics Li et al. (2024b).
4.4 Natural Science Research
Research has largely focused on applying LLMs to autonomous research workflows for natural science discovery. Auto-Bench Chen et al. (2025a) evaluated LLMs on chemistry and social science tasks based on causal graph discovery, revealing that LLMs perform effectively only when task complexity is highly limited. In contrast, ScienceAgentBench Chen et al. (2025b) provided a multidisciplinary benchmark for LLMs operating within agent frameworks such as CodeAct Wang et al. (2024b) and self-debug Chen et al. (2023). This benchmark highlighted the necessity for tailored agent workflows for such explorative tasks.
In the biomedical domain, Gao et al. (2024) discussed potential applications of AI agents in brainstorming, experimental planning, and execution. BioResearcher Luo et al. (2024) proposed an end-to-end framework for biomedical research involving dry lab experiments. DrugAgent Liu et al. (2025b) adopted multi-agent collaboration to automate drug discovery. In chemistry, Coscientist Boiko et al. (2023) incorporated the use of tools by LLMs to support semi-autonomous chemistry experiment design and execution. ProtAgents Ghafarollahi and Buehler (2024a) facilitated biochemistry discovery by building a multi-agent framework for automating protein design. Recent works, such as FutureHouse Skarlinski et al. (2025) and AI Co-scientist Gottweis et al. (2025), contributed to formulating demonstrably novel research hypotheses and proposals using multi-agent systems guided by predefined research goals.
4.5 General Research
Apart from specialized domain applications, some benchmarks have broadly evaluated diverse tasks from different stages of scientific discovery. DiscoveryWorld Jansen et al. (2024) created a virtual environment for LLM agents to conduct simplified scientific exploration. In Liu et al. (2025a), various application scenarios for AI agents in research were comprehensively discussed, supported by preliminary evaluation datasets. CURIE Cui et al. (2025) provides a comprehensive benchmark covering tasks in long-context understanding, reasoning, and information extraction, all of which are essential for scientific research. While EAIRA Cappello et al. (2025) focused on assessing the ability of LLMs to perform in a real-world research assistant role using various task formats.
5 Level 3. LLM as Scientist (Table A3)
Recently, several research efforts and commercial products have demonstrated prototypes of fully autonomous research within the artificial intelligence domain. These systems typically encompass a comprehensive workflow, from initial literature review to iterative refinement cycles where hypotheses or designs are progressively improved. A common feature is using an agent-based framework to autonomously produce research outputs, often culminating in draft research papers. This section will further compare these approaches, focusing on their methodologies for idea development and iterative refinement, as these aspects critically distinguish them from Level 2 agents.
5.1 Idea Development
The genesis of research in Level 3 systems involves transforming initial concepts into validated hypotheses, with distinct approaches to sourcing and vetting these ideas. Agent Laboratory Schmidgall et al. (2025) predominantly conducts literature reviews based on human-defined research objectives. Moving towards greater autonomy, several systems initiate their process from broader human inputs, such as reference papers Data Intelligence Lab (2025); Autoscience (2025) or general research domains IntologyAI (2025), subsequently exploring literature to autonomously identify gaps and formulate novel hypotheses. The AI Scientist (v1 Lu et al. (2024) and v2 Yamada et al. (2025)) showcases an even more generative approach: v1 brainstorms ideas from templates and past work, while v2 can generate diverse research proposals from abstract thematic prompts. Crucially, these systems employ diverse methods to evaluate their ideas prior to full implementation. AI Scientist-v1 uses self-assessed scores for interestingness, novelty, and feasibility, supplemented by external checks with Semantic Scholar. AI Scientist-v2 integrates literature review tools early in its idea formulation stage to assess novelty. This spectrum reveals a clear trend: while humans often initiate ideas, advanced systems can autonomously explore, generate, and validate the scientific merit and originality of research objectives before development.
5.2 Iterative Refinement
Iterative refinement within Level 3 systems involves sophisticated feedback loops that enable not just incremental improvements but also fundamental reassessments of the research trajectory. A key differentiator lies in the primary source and nature of this high-level feedback. The AI Scientist (v1 and v2) incorporates highly automated internal review and refinement processes. It employs AI reviewers, LLM evaluators for experimental choices, and VLMs to critique figures, fostering a rich internal feedback loop for iterative development. In contrast, Zochi IntologyAI (2025) integrates human expertise for macro-level guidance, where feedback can trigger complete re-evaluations of hypotheses or designs. This allows it to act on critiques challenging the core research premise, even reverting to hypothesis regeneration if results are unsatisfactory. Overall, while automated self-correction is a common goal, the current landscape reveals a pragmatic blend: some systems focus on enhancing autonomous deep reflection, while others integrate human oversight for robust, high-level iterative refinement and strategic redirection.
6 Challenges and Future Directions
Throughout this survey, we have systematically reviewed the escalating roles of Large Language Models in scientific discovery, delineating their progression through distinct levels of autonomy and capability—from foundational assistants and analysts to increasingly autonomous scientific researchers. In particular, we have underscored the evolving methodologies, task complexities, and the nature of human-LLM interaction that define each stage of this maturation. Beyond reviewing these advancements and current applications, this section presents several significant challenges and outlines promising directions for future research, aiming to inspire further exploration into the development and responsible deployment of LLMs as transformative tools in scientific inquiry.
Fully-Autonomous Research Cycle
While current Level 3 systems can navigate multiple stages of the scientific method for a specific inquiry, they often operate within a single research instance or predefined topic. The scientific method, however, is inherently cyclical, characterized by continuous iteration, refinement, and the pursuit of evolving research questions. A significant future direction, therefore, is to develop LLM-based systems capable of engaging in a truly autonomous research cycle. This would entail not merely executing a given research task from start to finish, but possessing the foresight to discern the broader implications of their findings, proactively identify promising avenues for subsequent investigation, and strategically direct their efforts towards practical advancements that build upon previous work.
Robotic Automation
A key barrier to fully autonomous scientific discovery in natural science is LLM agents’ inability to conduct physical laboratory experiments. While adept in computational research, their application in fields requiring physical interaction remains limited. Integrating LLMs with robotic systems empowers them to translate their planning capabilities into direct experimental actions. Early works in LLM-robotic integration Yoshikawa et al. (2023); Song et al. (2024); Darvish et al. (2025) already highlights this potential in chemical experimentation. Such automation is poised to significantly broaden LLM-based research, enabling end-to-end discovery in disciplines like chemistry and materials science, thereby advancing autonomous scientific exploration.
Transparency and Interpretability
The black-box nature (or opacity) of advancing LLMs in science undermines scientific validation, trust, and the assimilation of AI-driven insights Xu et al. (2025c). Addressing this opacity demands more than superficial Explainable AI (XAI) techniques Ahadian and Guan (2024). It necessitates a paradigm shift towards systems whose internal operations are inherently designed for verifiable reasoning and justifiable conclusions Bengio et al. (2025). Consequently, the challenge is not just explaining outputs, but ensuring the AI’s internal logic aligns with scientific principles and can reliably differentiate asserted claims from verifiable truths. This profound interpretability is vital for reliable and reproducible LLM-based scientific discovery.
Continuous Self-Improvement
The iterative and evolving nature of scientific inquiry demands systems capable of learning from ongoing engagement, assimilating experimental outcomes, and adapting research strategies. Current research integrating continual learning with agent-based systems already highlights the potential for LLMs to adapt to new tasks or environments without catastrophic forgetting Majumder et al. (2023); Kim et al. (2024). Within scientific discovery, a promising future direction is to incorporate online reinforcement learning frameworks Carta et al. (2023). This integration promises to continuously enhance scientific agents’ capabilities over their operational lifetime through successive discoveries, thereby advancing sustainable autonomous exploration.
Ethics and Societal Alignment
As LLM-based systems gain independent reasoning and action capabilities, their potential for risks—ranging from amplified societal biases to deliberate misuse like generating harmful substances or challenging human control—becomes increasingly salient and complex He et al. (2023); Ahadian and Guan (2024); Bengio et al. (2025). With AI capabilities and societal norms in constant flux, alignment is consequently an imperative, continuous process demanding adaptive governance and evolving value systems Li et al. (2024c). This requires embedding ethical constraints directly in scientific AI design frameworks, alongside vigilant oversight, to ensure advancements serve human well-being and the common good.
Limitations
This survey provides a systematic review of LLMs in scientific discovery, with a particular emphasis on the paradigm shift characterized by their escalating levels of autonomy. Our analysis and the selection of reviewed literature are therefore centered on works that illustrate this transition across the stages of the scientific method, categorized within our proposed three-level autonomy framework: LLM as Tool, LLM as Analyst, and LLM as Scientist.
Consequently, the scope of this survey has certain limitations. Firstly, we do not provide an exhaustive review of research focused on the development or fine-tuning of general-purpose scientific LLMs primarily designed for domain-specific knowledge acquisition, reasoning, or application. These areas, while crucial to the broader landscape of AI in science, are extensively covered in other existing surveys and fall outside our specific focus on the autonomy paradigm. Secondly, while we acknowledge the importance of fundamental LLM capabilities such as planning, code generation, and agentic decision-making, this survey does not delve deeply into orthogonal benchmarks or methodologies related to these general abilities unless they are explicitly contextualized and applied within the framework of LLM-driven scientific discovery and its evolving autonomy. These exclusions were deliberate to maintain a focused exploration of the transformative roles and increasing independence of LLMs throughout the scientific research lifecycle, rather than to offer a comprehensive catalog of all LLM-related tools or foundational capabilities that might be beneficial to scientific endeavors.
Ethics Statement
Our paper presents a comprehensive survey of LLMs in scientific discovery, with a specific focus on their role transformation from task automation tools to autonomous agents. All research works reviewed in this survey are properly cited. To the best of our knowledge, the referenced materials are publicly accessible or available under licenses permitting their use for research review. We did not conduct additional dataset curation or human annotation work. Consequently, we believe that this paper does not raise any ethical concerns.
Appendix A Summary Tables of LLMs in Scientific Discovery
Research Works Science Domain Task Nature
Framework
Methodology
Evaluation
Benchmark
Agentic
Workflow
Literature Search and Info Aggregation
LitLLM Agarwal et al. (2024) General Literature
Science Hierarchography Gao et al. (2025) General Literature
Dennstädt et al. (2024) Biomedicine Literature
SCIMON Wang et al. (2024a) General Literature, Idea Generation
ResearchAgent Baek et al. (2025) General Literature, Idea Generation
Text-Tuple-Table Deng et al. (2024b) General Text2Table
TKGT Jiang et al. (2024b) General Text2Table
ArxivDIGESTables Newman et al. (2024) General Literature, Text2Table
arXiv2Table Wang et al. (2025) General Literature, Text2Table
PaperQA & LitQA Lála et al. (2023) General Literature
Idea Generation and Hypothesis Formulation
Si et al. (2024) Artificial Intelligence Idea Generation
LiveIdeaBench Ruan et al. (2025) General Idea Generation
Nova Hu et al. (2024a) General Literature, Idea Generation
IdeaBench Guo et al. (2024a) General Literature, Idea Generation
GraphEval Feng et al. (2025) Artificial Intelligence Literature, Idea Generation
AI Idea Bench 2025 Qiu et al. (2025) Artificial Intelligence Literature, Idea Generation
Buehler (2024) Biology Literature, Idea Generation
SciAgents Ghafarollahi and Buehler (2024b) General Literature, Idea / Hypothesis Generation
MOOSE-Chem Yang et al. (2025) Chemistry Literature, Idea / Hypothesis Generation
Yang et al. (2024) General Literature, Idea / Hypothesis Generation
ResearchBench Liu et al. (2025c) General Literature, Idea / Hypothesis Generation
KG-CoI Xiong et al. (2024) General Literature, Idea / Hypothesis Generation
Ciucă et al. (2023) Astronomy Hypothesis Generation
O’Brien et al. (2024) Biomedicine Hypothesis Generation
LLM4GRN Afonja et al. (2024) Biology Hypothesis Generation
Zhou et al. (2024b) General Hypothesis Generation
Qi et al. (2023) General Hypothesis Generation
Qi et al. (2024) Biomedicine Hypothesis Generation
Scideator Radensky et al. (2025) General Idea / Hypothesis Generation
Experiment Planning and Execution
Li et al. (2025) General Planning
Shi et al. (2025) Biology Planning
BioPlanner O’Donoghue et al. (2023) Biology Planning
ARCADE Yin et al. (2022) Artificial Intelligence Code Generation
AIDE Jiang et al. (2025) Artificial Intelligence Code Generation
SciCode Tian et al. (2024) Artificial Intelligence Code Generation
DS-1000 Lai et al. (2022) Artificial Intelligence Code Generation
MLE-Bench Chan et al. (2025) Artificial Intelligence Code Generation
Data Analysis and Organization
AutomaTikZ Belouadi et al. (2024) General Text2Chart
Text2Chart31 Zadeh et al. (2025) General Text2Chart
ChartX & ChartVLM Xia et al. (2025) General Chart Reasoning
CharXiv Wang et al. (2024d) General Chart Reasoning
ChartQA Masry et al. (2022) General Chart Reasoning
Chain-of-Table Wang et al. (2024c) General Tabular Reasoning
TableBench Wu et al. (2025) General Tabular Reasoning
Deng et al. (2024a) General Tabular Reasoning
Conclusion and Hypothesis Validation
Tyser et al. (2024) General Review
ClaimCheck Ou et al. (2025) Artificial Intelligence Review
Du et al. (2024) Artificial Intelligence Review
Zhou et al. (2024a) Artificial Intelligence Review
ReviewerGPT Liu and Shah (2023) Artificial Intelligence Review
CycleResearcher Weng et al. (2025) Artificial Intelligence Review
Takagi et al. (2023) General Hypothesis Validation
Xu et al. (2025c) Physics Hypothesis Validation
SciReplicate-Bench Xiang et al. (2025) Artificial Intelligence Hypothesis Validation
Iteration and Refinement
Quan et al. (2024) General Refinement
MC-NEST Rabby et al. (2025) General Hypothesis Generation, Refinement
Chain of Ideas Li et al. (2024a) Artificial Intelligence Idea Generation, Refinement
Table A1: Comparison and classification of Level 1 research works in LLM-based scientific discovery.
Research Works Science Domain Methodology Framework Benchmark Evaluation Scientific Method Stages
Obs. Hyp. Exp. Ana. Con. Ref.
Machine Learning Research
CodeScientist Jansen et al. (2025) Artificial Intelligence
BudgetMLAgent Gandhi et al. (2025) Artificial Intelligence
IMPROVE Xue et al. (2025) Artificial Intelligence
MLAgentBench Huang et al. (2024a) Artificial Intelligence
Data Modeling and Analysis
DAgent Xu et al. (2025b) Data Science
DS-Agent Guo et al. (2024b) Data Science
InfiAgent-DABench Hu et al. (2024b) Data Science
BLADE Gu et al. (2024) Data Science
DiscoveryBench Majumder et al. (2024) Data Science
Zheng et al. (2023) General
Function Discovery
BoxLM Li et al. (2024b) Statistics
LLM-SR Shojaee et al. (2025a) Physics, Material, Biology
LLM-SRBench Shojaee et al. (2025b) Physics, Biology, Chemistry
Gravity-Bench-v1 Koblischke et al. (2025) Physics
Natural Science Research
Coscientist Boiko et al. (2023) Chemistry
Gao et al. (2024) Biomedicine
BioResearcher Luo et al. (2024) Biomedicine
DrugAgent Liu et al. (2025b) Biomedicine
FutureHouse Skarlinski et al. (2025) Chemistry, Biology
ScienceAgentBench Chen et al. (2025b) Chemistry, Biology
ProtAgents Ghafarollahi and Buehler (2024a) Chemistry, Biology
Auto-Bench Chen et al. (2025a) General
AI co-scientist Gottweis et al. (2025) General
General Research
DiscoveryWorld Jansen et al. (2024) General
Liu et al. (2025a) General
CURIE Cui et al. (2025) General
EAIRA Cappello et al. (2025) General
Table A2: Comparison and classification of Level 2 research works in LLM-based scientific discovery.
Research Works Science Domain Methodology Framework Benchmark Evaluation Featured Functionality Open-Sourced?
Agent Laboratory Schmidgall et al. (2025) Artificial Intelligence
literature review, experimentation, report writing,
iterative research with human feedback loops.
The AI Scientist Lu et al. (2024) Artificial Intelligence
idea generation, code generation,
experiment execution, research paper writing.
The AI Scientist-v2 Yamada et al. (2025) Artificial Intelligence
idea generation, code generation,
experiment execution, research paper writing,
with agentic tree-search and feedbacks.
AI-Researcher Data Intelligence Lab (2025) Artificial Intelligence
literature review, data analysis,
report generation.
Zochi IntologyAI (2025) Artificial Intelligence
customizable workflows for data collection,
analysis, and decision-making.
Carl Autoscience (2025) Artificial Intelligence
hypothesis generation, experiment design,
data analysis, and manuscript writing.
Table A3: Comparison and classification of Level 3 research works in LLM-based scientific discovery.